# The Very High Energy Afterglow of Structured Jets: GW 170817 and Prospects for Future Detections

### v1

@@ 2.4.5 The SSC case in Thomson regime
As $I_{2}/I_{0}$ is a function of $\gamma_{\mathrm{c}}/\gamma_{\mathrm{m}}=\gamma_{\mathrm{c}}^{\mathrm{syn}}/\gamma_{\mathrm{m}}/(1+Y^{\mathrm{no\,KN}})$, this is an implicit equation to be solved numerically to obtain $Y^{\mathrm{no\,KN}}$ and $\gamma_{\mathrm{c}}$. The limits for $\left(1+Y^{\mathrm{no\,KN}}\right)Y^{\mathrm{no\,KN}}$ are $\frac{3}{4}\frac{K_{\mathrm{IC}}}{K_{P}}\,\frac{\epsilon_{\mathrm{e}}}{\epsilon_{\mathrm{B}}}$ for $\gamma_{\mathrm{m}}\gg\gamma_{\mathrm{c}}$ and $\frac{3}{4}\frac{K_{\mathrm{IC}}}{K_{P}}\,\frac{\epsilon_{\mathrm{e}}}{\epsilon_{\mathrm{B}}}\frac{1}{3-p}\left(\frac{\gamma_{\mathrm{m}}}{\gamma_{\mathrm{c}}^{\mathrm{syn}}}\right)^{p-2}$ for $\gamma_{\mathrm{m}}\ll\gamma_{\mathrm{c}}$. Calculations using this prescription are labeled as ”SSC (Thomson)” in the following.
@@
Normalized Compton parameter.
In the general case where the KN suppression at high energy is included, the Compton parameter $Y(\gamma)$ is a decreasing function of the electron Lorentz factor. From Equations (40), (53) and (58), we get
@@
Note that for low Lorentz factors $\gamma$, $\widetilde{\nu}$ becomes very large so that, for $\widetilde{\nu}>\nu_{\mathrm{syn}}(\gamma_{\mathrm{max}})$, we recover the Thomson regime, where the right-hand side of Equation (78) is formally the same as in Equation (77), even if the integrals $I_{0}$ and $I_{2}$ can be different if the normalized electron distribution $f(x)$ is different. As described in Nakar et al. (2009), when keeping only the leading term in the integral of $g(y)$, the corresponding scaling law for the Compton parameter is $Y(\gamma)=\mathrm{cst}$ if $\widetilde{\nu}$ is above the peak frequency $\nu_{\mathrm{p}}$ and $Y(\gamma)\propto\gamma^{-\frac{3-\widetilde{p}}{2}}$ otherwise, where $\widetilde{p}$ is the slope of electron distribution in the power-law segment of the electron distribution including $\widetilde{\gamma}$. This leads us to introduce a normalized Compton parameter defined by
@@
where $h(x)$ follows this scaling and is normalized so that Equation (79) has the exact limit in the Thomson regime. This leads to
$\!\!\!h(x)\!=\!\left\{\begin{array}[]{rl}1&\!\mathrm{if}\,x\!<\widehat{x}_{\mathrm{p}}\\ \widehat{x}_{\mathrm{p}}^{\frac{3-p_{i_{\mathrm{p}}-1}}{2}}x^{-\frac{3-p_{i_{\mathrm{p}}-1}}{2}}&\!\mathrm{if}\,\widehat{x}_{\mathrm{p}}<\!x\!<\widehat{x}_{i_{\mathrm{p}}-1}\\ \!\!\!\widehat{x}_{\mathrm{p}}^{\frac{3-p_{i_{\mathrm{p}}-1}}{2}}\widehat{x}_{i_{\mathrm{p}}-1}^{\frac{p_{i_{\mathrm{p}}-1}-p_{i_{\mathrm{p}}-2}}{2}}x^{-\frac{3-p_{i_{\mathrm{p}}-2}}{2}}&\!\mathrm{if}\,\widehat{x}_{i_{\mathrm{p}}-1}<\!x\!<\widehat{x}_{i_{\mathrm{p}}-2}\\ \vdots&\vdots\\ \widehat{x}_{\mathrm{p}}^{\frac{3-p_{i_{\mathrm{p}}-1}}{2}}\widehat{x}_{i_{\mathrm{p}}-1}^{\frac{p_{i_{\mathrm{p}}-1}-p_{i_{\mathrm{p}}-2}}{2}}\cdots&\\ \cdots\,\widehat{x}_{2}^{\frac{p_{2}-p_{1}}{2}}x^{-\frac{3-p_{1}}{2}}&\!\mathrm{if}\,\widehat{x}_{2}<\!x<\widehat{x}_{1}\\ \widehat{x}_{\mathrm{p}}^{\frac{3-p_{i_{\mathrm{p}}-1}}{2}}\widehat{x}_{i_{\mathrm{p}}-1}^{\frac{p_{i_{\mathrm{p}}-1}-p_{i_{\mathrm{p}}-2}}{2}}\cdots&\\ \cdots\,\!\widehat{x}_{2}^{\frac{p_{2}-p_{1}}{2}}\widehat{x}_{1}^{\frac{p_{1}}{2}-\frac{1}{6}}x^{-4/3}&\!\mathrm{if}\,x\!>\widehat{x}_{1}\end{array}\right.$ (80)
@@
Self-consistent solution for the normalized electron distribution.
Following Nakar et al. (2009), we define [^2] As $Y(\gamma)$ is a decreasing function with a constant first segment, if $Y(\gamma)<1$ for $\gamma<\max{\left(\gamma_{\mathrm{c}};\widehat{\gamma}_{\mathrm{p}}\right)}$, then $Y(\gamma)<1$ for all values of $\gamma$, and $\gamma_{0}$ is therefore undefined. In these cases, we retrieve the ”no SSC” solution. $\gamma_{0}$ by $Y(\gamma_{0})=1$ and assume $1+Y(\gamma)=Y(\gamma)$ for $\gamma<\gamma_{0}$ and $1$ otherwise. From Equations (44) and (LABEL:eq:barnSC), this shows that the IC cooling will only affect the electron distribution in the interval
@@
Therefore the solution is entirely determined by the orderings of $\gamma_{\mathrm{m}}$, $\gamma_{\mathrm{c}}$, $\gamma_{0}$, $\widehat{\gamma}_{\mathrm{m}}$ and $\widehat{\gamma_{\mathrm{c}}}$. All possible cases and the corresponding solutions are listed in Appendix B. In some cases, subcases are introduced as breaks can appear at $\widehat{\gamma}_{0}$, $\widehat{\widehat{\gamma}}_{\mathrm{m}}$, $\widehat{\widehat{\gamma}}_{\mathrm{c}}$, etc. All the most relevant cases for GRB afterglows are already described in Nakar et al. (2009) along with the detailed method to obtain the corresponding solution for the electron distribution. For completeness, we list in Appendix B the additional cases allowing to fully describe the parameter space, even in regions unlikely to be explored in GRBs.
@@
From Equation (79), we get the following equation for $Y(\gamma_{\mathrm{c}})$:
$Y(\gamma_{\mathrm{c}})=\left\{\begin{array}[]{cl}A\,\frac{\gamma_{\mathrm{m}}}{\gamma_{\mathrm{c}}^{\mathrm{syn}}}\,\frac{I_{2}}{I_{0}}\,h(x_{\mathrm{c}})&\mathrm{if}\,\gamma_{\mathrm{m}}>\gamma_{\mathrm{c}}\\ A\,\frac{\gamma_{\mathrm{c}}^{\mathrm{syn}}}{\gamma_{\mathrm{m}}}\,\frac{I_{2}}{I_{0}}\,h(x_{\mathrm{c}})&\mathrm{if}\,\gamma_{\mathrm{m}}<\gamma_{\mathrm{c}}\,\mathrm{and}\,Y(\gamma_{\mathrm{c}})<1\\ \left(A\,\frac{\gamma_{\mathrm{c}}^{\mathrm{syn}}}{\gamma_{\mathrm{m}}}\,\frac{I_{2}}{I_{0}}\,h(x_{\mathrm{c}})\right)^{1/3}&\mathrm{if}\,\gamma_{\mathrm{m}}<\gamma_{\mathrm{c}}\,\mathrm{and}\,Y(\gamma_{\mathrm{c}})>1\\ \end{array}\right.$ (83)
where $A=\frac{3}{4}\frac{K_{\mathrm{IC}}}{K_{P}}\,\frac{p-2}{p-1}\,\frac{\epsilon_{\mathrm{e}}}{\epsilon_{\mathrm{B}}}$ is a constant as the microphysics parameter $\epsilon_{\mathrm{B}}$, $\epsilon_{\mathrm{e}}$ and $p$ are assumed to be constant during the forward shock propagation. Note that $x_{\mathrm{c}}=1$ in the slow cooling regime $\gamma_{\mathrm{m}}<\gamma_{\mathrm{c}}$. A second useful relation is obtained from Equation (79) and the definition of $\gamma_{0}$:
@@
Then, $\gamma_{\mathrm{c}}$ is computed from the following implicit equation, which is solved iteratively:
@@
We start by assuming $Y(\gamma_{\mathrm{c}})=Y^{\mathrm{no\,KN}}$ as defined in § 2.4.5 and then iterate as follows:
@@
    Compute $\gamma_{\mathrm{c}}$ from Equation (85).
@@
    Knowing $\gamma_{\mathrm{m}}$, $\gamma_{\mathrm{c}}$, $\gamma_{\mathrm{self}}$, scan the different possible cases for the ordering of $\gamma_{\mathrm{m}}$, $\gamma_{\mathrm{c}}$, $\gamma_{0}$, $\widehat{\gamma}_{\mathrm{m}}$ and $\widehat{\gamma_{\mathrm{c}}}$ (and additional characteristic Lorentz factors for some cases) as listed in Appendix B and compute $\gamma_{0}$ from Equation (84), which can be analytically inverted as provided for each case in Appendix B. We stop the scan of the possible cases when the ordering of all characteristic Lorentz factors including the obtained value of $\gamma_{0}$ is the correct one.
@@
    Compute an updated value of $Y(\gamma_{\mathrm{c}})$ from Equation (83).
@@
We stop the procedure when the relative variation of $\gamma_{\mathrm{c}}$ in an iteration falls below $\epsilon_{\mathrm{tol}}$. We use $\epsilon_{\mathrm{tol}}=10^{-4}$, which is reached in most cases in less than $20$ iterations. In the context of a full afterglow light curve calculation, as we expect $Y$ to vary smoothly during the jet propagation, we use the value of $Y(\gamma_{\mathrm{c}})$ obtained at the previous step of the dynamics instead of $Y^{\mathrm{no\,KN}}$ to start the iterative procedure. Then we usually reach a convergence in only a few iterations.
@@ 2.4.7 Final calculation of the emission in the comoving frame
The self-consistent spectrum in the comoving frame, i.e. $p_{\nu}^{\mathrm{syn}}$ and $p_{\nu}^{\mathrm{SSC}}$, the synchrotron and SSC powers per electron and per unit frequency averaged over the timescale $t_{\mathrm{dyn}}$, can be computed by the following procedure based on the previous paragraphs:
@@
    The following quantities can be immediately deduced: $\gamma_{\mathrm{c}}^{\mathrm{syn}}$ from Equation (43), $\gamma_{\mathrm{self}}$ from Equation (38), $\tau_{\mathrm{T}}$ from Equation (61), and $Y^{\mathrm{no\,KN}}$ from Equation (77).
@@
    Then the iterative procedure described above allows to identify the spectral regime among all the possibilities listed in Appendix B, and to compute $\gamma_{\mathrm{c}}$. At the end of this step, all breaks and slopes in the functions $f(x)$ and $h(x)$ are known. The corresponding integrals $I_{0}$ and $I_{2}$ are computed by a numerical integration of Equations (51) and (55).
@@
    The function $g(y)$ can immediately be deduced from $f(x)$, using Equation (54). The corresponding integral $J_{0}$ is computed by a numerical integration of Equation (56), and the synchrotron spectrum per electron, $p_{\nu}^{\mathrm{syn}}$, is deduced from Equation (53).
@@
    Finally, the function $G(y)$ is computed by an numerical integration of Equation (62) and the SSC spectrum per electron, $p_{\nu}^{\mathrm{SSC}}$, is deduced from Equation (60).

### v2

@@ 2.4.5 Synchrotron self-Compton case in the Thomson regime
As $I_{2}/I_{0}$ is a function of $\gamma_{\mathrm{c}}/\gamma_{\mathrm{m}}=\gamma_{\mathrm{c}}^{\mathrm{syn}}/\gamma_{\mathrm{m}}/(1+Y^{\mathrm{no\,KN}})$, this is an implicit equation to be solved numerically to obtain $Y^{\mathrm{no\,KN}}$ and $\gamma_{\mathrm{c}}$. The limits for $\left(1+Y^{\mathrm{no\,KN}}\right)Y^{\mathrm{no\,KN}}$ are $\frac{3}{4}\frac{K_{\mathrm{IC}}}{K_{P}}\,\frac{\epsilon_{\mathrm{e}}}{\epsilon_{\mathrm{B}}}$ for $\gamma_{\mathrm{m}}\gg\gamma_{\mathrm{c}}$ and $\frac{3}{4}\frac{K_{\mathrm{IC}}}{K_{P}}\,\frac{\epsilon_{\mathrm{e}}}{\epsilon_{\mathrm{B}}}\frac{1}{3-p}\left(\frac{\gamma_{\mathrm{m}}}{\gamma_{\mathrm{c}}^{\mathrm{syn}}}\right)^{p-2}$ for $\gamma_{\mathrm{m}}\ll\gamma_{\mathrm{c}}$. Calculations using this prescription are labelled ”SSC (Thomson)” in the following.
@@
Normalised Compton parameter.
In the general case where the KN suppression at high energy is included, the Compton parameter $Y(\gamma)$ is a decreasing function of the electron Lorentz factor. From Eqs. (40), (53), and (58), we obtain
@@
For low Lorentz factors $\gamma$, $\widetilde{\nu}$ becomes very large, so that for $\widetilde{\nu}>\nu_{\mathrm{syn}}(\gamma_{\mathrm{max}})$, we recover the Thomson regime, where the right-hand side of Eq. (78) is formally the same as in Eq. (77), even if the integrals $I_{0}$ and $I_{2}$ can be different when the normalised electron distribution $f(x)$ is different. As described in Nakar et al. (2009), when only the leading term in the integral of $g(y)$ is kept, the corresponding scaling law for the Compton parameter is $Y(\gamma)=\mathrm{cst}$ if $\widetilde{\nu}$ is above the peak frequency $\nu_{\mathrm{p}}$ and $Y(\gamma)\propto\gamma^{-\frac{3-\widetilde{p}}{2}}$ otherwise, where $\widetilde{p}$ is the slope of electron distribution in the power-law segment of the electron distribution including $\widetilde{\gamma}$. We therefore introduce a normalised Compton parameter defined by
@@
where $h(x)$ follows this scaling and is normalised so that Eq. (79) has the exact limit in the Thomson regime. This leads to
$h(x)=\left\{\begin{array}[]{rl}1&\!\mathrm{if}\,x\!<\widehat{x}_{\mathrm{p}}\\ \widehat{x}_{\mathrm{p}}^{\frac{3-p_{i_{\mathrm{p}}-1}}{2}}x^{-\frac{3-p_{i_{\mathrm{p}}-1}}{2}}&\!\mathrm{if}\,\widehat{x}_{\mathrm{p}}<\!x\!<\widehat{x}_{i_{\mathrm{p}}-1}\\ \!\!\!\widehat{x}_{\mathrm{p}}^{\frac{3-p_{i_{\mathrm{p}}-1}}{2}}\widehat{x}_{i_{\mathrm{p}}-1}^{\frac{p_{i_{\mathrm{p}}-1}-p_{i_{\mathrm{p}}-2}}{2}}x^{-\frac{3-p_{i_{\mathrm{p}}-2}}{2}}&\!\mathrm{if}\,\widehat{x}_{i_{\mathrm{p}}-1}<\!x\!<\widehat{x}_{i_{\mathrm{p}}-2}\\ \vdots&\vdots\\ \widehat{x}_{\mathrm{p}}^{\frac{3-p_{i_{\mathrm{p}}-1}}{2}}\widehat{x}_{i_{\mathrm{p}}-1}^{\frac{p_{i_{\mathrm{p}}-1}-p_{i_{\mathrm{p}}-2}}{2}}\cdots&\\ \cdots\,\widehat{x}_{2}^{\frac{p_{2}-p_{1}}{2}}x^{-\frac{3-p_{1}}{2}}&\!\mathrm{if}\,\widehat{x}_{2}<\!x<\widehat{x}_{1}\\ \widehat{x}_{\mathrm{p}}^{\frac{3-p_{i_{\mathrm{p}}-1}}{2}}\widehat{x}_{i_{\mathrm{p}}-1}^{\frac{p_{i_{\mathrm{p}}-1}-p_{i_{\mathrm{p}}-2}}{2}}\cdots&\\ \cdots\,\!\widehat{x}_{2}^{\frac{p_{2}-p_{1}}{2}}\widehat{x}_{1}^{\frac{p_{1}}{2}-\frac{1}{6}}x^{-4/3}&\!\mathrm{if}\,x\!>\widehat{x}_{1}\end{array}\right..$ (80)
@@
Self-consistent solution for the normalised electron distribution.
Following Nakar et al. (2009), we define [^2] As $Y(\gamma)$ is a decreasing function with a constant first segment, if $Y(\gamma)<1$ for $\gamma<\max{\left(\gamma_{\mathrm{c}};\widehat{\gamma}_{\mathrm{p}}\right)}$, then $Y(\gamma)<1$ for all values of $\gamma$, and $\gamma_{0}$ is therefore undefined. In these cases, we retrieve the ”no SSC” solution. $\gamma_{0}$ by $Y(\gamma_{0})=1$ and assume $1+Y(\gamma)=Y(\gamma)$ for $\gamma<\gamma_{0}$ and $1$ otherwise. From Eqs. (44) and (LABEL:eq:barnSC), this shows that the IC cooling only affects the electron distribution in the interval
@@
Therefore, the solution is entirely determined by the orderings of $\gamma_{\mathrm{m}}$, $\gamma_{\mathrm{c}}$, $\gamma_{0}$, $\widehat{\gamma}_{\mathrm{m}}$, and $\widehat{\gamma_{\mathrm{c}}}$. All possible cases and the corresponding solutions are listed in Appendix B. In some cases, subcases are introduced as breaks can appear at $\widehat{\gamma}_{0}$, $\widehat{\widehat{\gamma}}_{\mathrm{m}}$, $\widehat{\widehat{\gamma}}_{\mathrm{c}}$, etc. All the most relevant cases for GRB afterglows were described in Nakar et al. (2009), along with the detailed method for obtaining the corresponding solution for the electron distribution. For completeness, we list in Appendix B the additional cases that allow us to fully describe the parameter space, even in regions that are unlikely to be explored in GRBs.
@@
From Eq. (79), we obtain the following equation for $Y(\gamma_{\mathrm{c}})$:
$Y(\gamma_{\mathrm{c}})=\left\{\begin{array}[]{cl}A\,\frac{\gamma_{\mathrm{m}}}{\gamma_{\mathrm{c}}^{\mathrm{syn}}}\,\frac{I_{2}}{I_{0}}\,h(x_{\mathrm{c}})&\mathrm{if}\,\gamma_{\mathrm{m}}>\gamma_{\mathrm{c}}\\ A\,\frac{\gamma_{\mathrm{c}}^{\mathrm{syn}}}{\gamma_{\mathrm{m}}}\,\frac{I_{2}}{I_{0}}\,h(x_{\mathrm{c}})&\mathrm{if}\,\gamma_{\mathrm{m}}<\gamma_{\mathrm{c}}\,\mathrm{and}\,Y(\gamma_{\mathrm{c}})<1\\ \left(A\,\frac{\gamma_{\mathrm{c}}^{\mathrm{syn}}}{\gamma_{\mathrm{m}}}\,\frac{I_{2}}{I_{0}}\,h(x_{\mathrm{c}})\right)^{1/3}&\mathrm{if}\,\gamma_{\mathrm{m}}<\gamma_{\mathrm{c}}\,\mathrm{and}\,Y(\gamma_{\mathrm{c}})>1\\ \end{array}\right.,$ (83)
where $A=\frac{3}{4}\frac{K_{\mathrm{IC}}}{K_{P}}\,\frac{p-2}{p-1}\,\frac{\epsilon_{\mathrm{e}}}{\epsilon_{\mathrm{B}}}$ is a constant as the microphysics parameters $\epsilon_{\mathrm{B}}$, $\epsilon_{\mathrm{e}}$, and $p$ are assumed to be constant during the forward-shock propagation. We note that $x_{\mathrm{c}}=1$ in the slow-cooling regime $\gamma_{\mathrm{m}}<\gamma_{\mathrm{c}}$. A second useful relation is obtained from Eq. (79) and the definition of $\gamma_{0}$,
@@
Then, $\gamma_{\mathrm{c}}$ is computed from the following implicit equation, which is solved iteratively,
@@
We start by assuming $Y(\gamma_{\mathrm{c}})=Y^{\mathrm{no\,KN}}$ as defined in Sect. 2.4.5 and then iterate as follows:
@@
    Compute $\gamma_{\mathrm{c}}$ from Eq. (85).
@@
    Knowing $\gamma_{\mathrm{m}}$, $\gamma_{\mathrm{c}}$, and $\gamma_{\mathrm{self}}$, scan the different possible cases for the ordering of $\gamma_{\mathrm{m}}$, $\gamma_{\mathrm{c}}$, $\gamma_{0}$, $\widehat{\gamma}_{\mathrm{m}}$, and $\widehat{\gamma_{\mathrm{c}}}$ (and additional characteristic Lorentz factors for some cases) as listed in Appendix B and compute $\gamma_{0}$ from Eq. (84), which can be analytically inverted as provided for each case in Appendix B. We stop the scan of the possible cases when the ordering of all characteristic Lorentz factors including the obtained value of $\gamma_{0}$ is the correct one.
@@
    Compute an updated value of $Y(\gamma_{\mathrm{c}})$ from Eq. (83).
@@
We stop the procedure when the relative variation of $\gamma_{\mathrm{c}}$ in an iteration falls below $\epsilon_{\mathrm{tol}}$. We use $\epsilon_{\mathrm{tol}}=10^{-4}$, which is reached in most cases in fewer than $20$ iterations. In the context of a full afterglow light-curve calculation, as we expect $Y$ to vary smoothly during the jet propagation, we use the value of $Y(\gamma_{\mathrm{c}})$ obtained at the previous step of the dynamics instead of $Y^{\mathrm{no\,KN}}$ to start the iterative procedure. Then, we usually reach convergence in only a few iterations.
@@ 2.4.7 Final calculation of the emission in the comoving frame
The self-consistent spectrum in the comoving frame, that is, $p_{\nu}^{\mathrm{syn}}$ and $p_{\nu}^{\mathrm{SSC}}$, the synchrotron and SSC powers per electron and per unit frequency averaged over the timescale $t_{\mathrm{dyn}}$, can be computed by the following procedure based on the previous paragraphs:
@@
    The following quantities can immediately be deduced: $\gamma_{\mathrm{c}}^{\mathrm{syn}}$ from Eq. (43), $\gamma_{\mathrm{self}}$ from Eq. (38), $\tau_{\mathrm{T}}$ from Eq. (61), and $Y^{\mathrm{no\,KN}}$ from Eq. (77).
@@
    Then, the iterative procedure described above allows us to identify the spectral regime among all the possibilities listed in Appendix B, and we can compute $\gamma_{\mathrm{c}}$. At the end of this step, all breaks and slopes in the functions $f(x)$ and $h(x)$ are known. The corresponding integrals $I_{0}$ and $I_{2}$ are computed by a numerical integration of Eqs. (51) and (55).
@@
    The function $g(y)$ can immediately be deduced from $f(x)$, using Eq. (54). The corresponding integral $J_{0}$ is computed by a numerical integration of Eq. (56), and the synchrotron spectrum per electron, $p_{\nu}^{\mathrm{syn}}$, is deduced from Eq. (53).
@@
    Finally, the function $G(y)$ is computed by an numerical integration of Eq. (62) and the SSC spectrum per electron, $p_{\nu}^{\mathrm{SSC}}$, is deduced from Eq. (60).